OpenStudy (anonymous):

Solve the quadratic equation by completing the square. 2x^2 + 10x - 2 = 0

OpenStudy (anonymous):

2x^2 + 10x - 2 = 0 a b c get rid of c put it on the right then divide by a to get rid of a 2x^2 + 10x = 2 x^2 + 5x = 1 ------- completing the square comes from a perfect square (x+a)*(x+a) when you distribute x^2+2ax+a^2 NOTE this is not the same a as above new a new variable. --------- wee need to add something to x^2 + 5x = 1 to make it a perfect square. x^2 +2ax +a^2 If we put the distrubets (x+a)^2 under it we see that 5=2a aka the coeeficient so if we solve for a a=5/2 further more we see that a has to be squared so we add a^2 to both sides x^2 + 5x +(5/2)^2 = 1+ (5/2)^2 now since x^2 + 5x +(5/2)^2 is a perfect square (x+a)^2 and we know a=5/2 we make left hand a perfect square by plugin in a (x+5/2)^2= 1+ (5/2)^2 we take square root now (x+5/2)= plus or minussqrt(1+ (5/2)^2) and solve normally now